### v2

# DeblurDiNAT: A Lightweight and Effective Transformer for Image Deblurring

Abstract
While recent Transformer-based architectures generate improved deblurring outcomes than prior convolutional neural networks (CNNs), the large model size and heavy computational cost are still two bothersome issues which have not been fully resolved. Besides, the existing state-of-the-art (SOTA) approach trained on synthetic datasets may not perform well on a broad range of unseen real-world deblurring scenarios. To this end, we propose DeblurDiNAT, a compact and robust encoder-decoder Transformer which efficiently restores clean images from blurred ones. We adopt an alternating dilation factor structure to learn local and global blur patterns. Also, we observe that Transformers with conventional self-attention layers do not always produce good deblurred results. To solve this problem, we propose a channel modulation self-attention (CMSA) block, where a local cross-channel learner (LCCL) is utilized to capture channel relationships. In addition, we present a divide and multiply feed-forward network (DMFN) allowing fast feature propagation. Moreover, we design a lightweight dual-stage feature fusion (LDFF) module to introduce non-linearity. Comprehensive experimental results show that the proposed DeblurDiNAT provides a significant performance boost without a noticeable increase in computational cost over the baseline, and achieves SOTA performance on mainstream image deblurring datasets. Compared to nearest competitors, our space-efficient and time-saving model demonstrates a stronger generalization ability with $3\%$-$68\%$ fewer parameters and produces deblurred images that are visually closer to the ground truth. Source codes and pre-trained models are available at https://github.com/HanzhouLiu/DeblurDiNAT.git.
1 Introduction
Image deblurring is the procedure of removing blur artifacts from unclear or blurred images [5, 43], and is traditionally conceptualized as an inverse filtering problem [2, 13, 31]. Non-deep learning methods exhibit robust efficacy in certain scenarios [4, 1, 18, 42]. However, their performance is suboptimal in more intricate yet prevalent cases, for example, those involving strong motion blur [49].
In the past decade, convolutional neural networks (CNN) play an important role in computer vision tasks [8, 48, 27]. The significant strength of CNNs is their ability to acquire generalizable image priors from large-scale datasets, making them become de facto for image deblurring [25, 37, 33, 6, 50, 22, 11, 29]. Nonetheless, the limited receptive field constrained by the kernel size and network depth prevents CNNs from modeling long-range pixel dependencies effectively.
Transformers, initially designed for machine translation [39], have adapted to handle visual data, challenging the dominance of CNNs in computer vision [10, 30, 44, 23]. Self-attention (SA), a fundamental component in Transformers, differs from convolutional layers in CNNs and facilitates long-range feature learning through scaled dot-product. This attribute may hinder the capacity of Transformers in capturing local blurred patterns that exist in blurred images [14, 13, 38]. Besides, computational cost increases quadratically with respect to the input size. To overcome these limitations, most methods split image data into small spatial windows or introduce conventional 2-D convolutions [41, 38, 45, 24]. Although the model complexity is reduced and short-range visual representations are learnt in this way, the critical issue of partial loss of both spatial and cross-channel information has not yet been resolved. The existing SOTA deblurring model [24] tested on synthetic datasets is not able to generalize well from the known samples to a wide range of unseen real-world scenarios. It has proven to be a challenging task for deblurring Transformers to achieve the optimal balance among effectiveness, efficiency, and robustness.
In this paper, we intend to construct a deblurring Transformer with a compact size, low computational cost and SOTA generalization ability. We carefully investigate recent Transformer work for the deblurring task [41, 38, 45, 24] and propose three questions. First, is there an efficient way to capture both local and global blur patterns commonly existing in blurred images? Second, prior Transformers benefit from 2-D convolutions to learn cross-channel relationships globally [45, 24]. Is there an lightweight approach for Transformers to understand local channel dependencies that play an important role in image data either? Third, a simple gate without any gating mechanism functions effectively within a CNN architecture [3]. Does that hold true for a deblurring Transformer? If so, how to ensure non-linearity in order to model complex visual data? To answer these questions, we propose DeblurDiNAT, an effective, efficient and robust deblurring Transformer, built on a novel hybrid encoder-decoder architecture.
As one of the core mechanisms of DeblurDiNAT, our SA strategy is based on Dilated Neighborhood Attention (DiNA) [15, 16], a flexible and powerful sparse global attention pattern. A dilation factor is used to adjust the receptive field size without additional overhead. A larger dilation factor results in a larger receptive field [21]. Through stacking Transformer blocks with small and large dilation factors alternately, our network learns both local and global blur information in an efficient way. To the best of our knowledge, this paper is the first work that comprehensively explores the potential of DiNA in an image restoration task.
To enable Transformers to understand local channel dependencies, we introduce a local cross-channel learner (LCCL) that only considers each channel’s nearest neighbors. LCCL draws inspiration from existing channel attention methods [20, 40]. Global channel relationships can be easily learnt by 2-D convolutions, which is not the focus of this paper. Formally, we propose a new SA framework, named channel modulation self-attention (CMSA), where a parallel LCCL dynamically modulates the outputs of an SA block (e.g., DiNA) based on its input feature maps. Although our parallel design was initially developed for efficiency, it outperforms other variants in terms of metric scores.
The other basic component of Transformers is the feed-forward network (FN). There are two common types of FNs in SOTA deblurring Transformers, FNs composed of fully connected layers followed by activations [38, 41], and the ones based on 2-D convolutions [45, 24]. In this paper, we aim to investigate and optimize the second type, where a GELU [17] function is frequently used for the purpose of non-linearity. Questions may arise regarding the efficiency of this approach, considering that Transformers are built on sequentially stacked FNs and SA blocks. To further reduce computational loads, we present a divide and multiply feed-forward network (DMFN) without any non-linear activation function. Note that the non-linearity is handled in other modules. Experimental results show that this clean design not only decreases inference time but also positively influences deblurring outcomes.
Apart from the above novelties, we visit a different component of the network, feature fusion. We make the following empirical observations. First, existing Transformers for image deblurring only uses same-scale feature fusion, which cannot fully compensate for the information loss caused by the downsampling process, especially in deeper levels. Second, prior feature fusion approaches [28, 7], originally designed for CNNs, add extra computational complexity to Transformers due to unneeded convolutions, activations and normalizations. Third, as no activation exists in the decoder path of DeblurDiNAT, non-linearity should be introduced somewhere to handle and propagate complex visual data across the network. Therefore, we propose a non-linear lightweight dual-stage feature fusion (LDFF) module, an efficient method to merge multi-scale or same-scale features.
Our contributions in this work are summarized as follows:
  • We propose DeblurDiNAT, an encoder-decoder Transformer for image deblurring, capable of capturing both local and global blurred patterns.
  • We develop a channel modulation self-attention (CMSA) block, allowing the Transformer to learn local cross-channel relationships in an efficient way.
  • We present a divide and multiply feed-forward network (DMFN), which propagates features with optimal computational cost.
  • We design a lightweight dual-stage feature fusion (LDFF) approach that is well-suited for both multi-scale and same-scale feature aggregation.
  • Extensive experiments show that DeblurDiNAT is an efficient deblurring Transformer that performs on par with SOTA models, and demonstrate a very strong generalization ability on mainstream datasets.
2 Background
CNN-based Image Deblurring Models. Deep CNN-based methods have proven to be highly effective in image deblurring [19, 36, 32, 26, 37, 12, 47, 6, 46]. Tao et al. [37] investigated the coarse-to-fine scheme, and proposed a Scale-recurrent Network (SRN) with a simpler structure compared to conventional cascaded neural networks. Cho et al. [6] revisited the coarse-to-fine strategy and presented a multi-input multi-output UNet (MIMO-UNet) for fast image deblurring. Zamir et al. [46] chose to utilize multi-level representations and designed a multi-stage progressive architecture (MPRNet) to significantly enhance the deblurring performance.
Transformer-based Image Deblurring Models. Researchers have dedicated significant efforts to capturing spatial variant blur, in order to improve the applicability of Transformers in real-world image deblurring [41, 38, 45, 24]. Tsai et al. [38] leveraged intra-strip and inter-strip attention for estimating various blurred patterns and constructed Stripformer for dynamic scene deblurring [38]. Zamir et al. [45] proposed Restormer, an efficient Transformer which computes cross-covariance across feature channels rather than spatial dimension to model global connectivity. Although these approaches have produced better deblurring results than prevalent CNN models, they have not fully utilized cross-channel relationships in complex multi-channel data. Kong et al. [24] advanced deblurring Transformers by introducing quantization and Fourier transform to the basic building blocks, which achieves the highest metric score tested on the synthetic GoPro dataset [32]. However, it demonstrates a very weak generalization ability to process unknown real-world blurred images.
Multi-scale Feature Fusion. Li et al. [28] presented a Gated Fully Fusion (GFF) mechanism using multiple addition gates and element-wise multiplications to generate high-level feature maps for semantic segmentation. Dong et al. [9] proposed a Dense Feature Fusion (DFF) module for U-Net architectures, based on the back-projection technique. Nevertheless, redundant matrix operations in GFF and stacked convolutional layers in DFF result in heavy computational burdens. Cho et al. [6] adopted Asymmetric Feature Fusion (AFF) composed of only two convolutional layers to allow fast information flow across a U-Net, which might be a limiting factor to learn complex representations.
3 Method
DeblurDiNAT is an efficient hybrid Transformer for single image deblurring. In this section, we first introduce DiNA [15] to image deblurring and present an alternating dilation factor structure to capture local and global blurred patterns. Then, we detail the channel modulation self-attention (CMSA), leveraging local cross-channel relationships to model complex multi-channel data. After that, we present divide and multiply feed-forward network (DMFN) for fast feature propagation. Lastly, we provide a lightweight dual-stage feature fusion (LDFF) module to further reduce computational cost and improve image quality. Fig. 1 shows the architecture of DeblurDiNAT. Each encoder performs convolutions followed by Leaky ReLU activations, where convolutions extract hierarchical features and Leaky ReLU contributes to a robust learning process. The other components are explained in the following.
[…]
$\mathbf{A}_{i}^{(k,\delta)}=\begin{bmatrix}Q_{i}K_{\rho_{1}^{\delta}(i)}^{T}+B_{(i,\rho_{1}^{\delta}(i))}\\ Q_{i}K_{\rho_{2}^{\delta}(i)}^{T}+B_{(i,\rho_{2}^{\delta}(i))}\\ \vdots\\ Q_{i}K_{\rho_{k}^{\delta}(i)}^{T}+B_{(i,\rho_{k}^{\delta}(i))}\end{bmatrix}.$ (1)
[…]
$\mathbf{V}_{i}^{(k,\delta)}=\begin{bmatrix}V_{\rho_{1}^{\delta}(i)}^{T}&V_{\rho_{2}^{\delta}(i)}^{T}&\cdots&V_{\rho_{k}^{\delta}(i)}^{T}\end{bmatrix}^{T}.$ (2)
[…]
$\mathrm{DiNA}_{k}^{\delta}(i)=\emph{softmax}\>(\frac{\mathbf{A}_{i}^{(k,\delta)}}{\sqrt{d_{k}}})\mathbf{V}_{i}^{(k,\delta)}.$ (3)
Figure 1: Architecture of DeblurDiNAT for single image deblurring. DeblurDiNAT is a hybrid encoder-decoder network, where cascaded residual blocks extract multi-scale image features from input images, and alternating local and global Transformer blocks leverage hierarchical features for image reconstruction. Lightweight dual-stage feature fusion (LDFF) modules perform effective and flexible feature fusion across the model.
We refer our readers to the original work for complexity analysis [16].
3.1 Alternating Dilation Factor Structure
In DiNA, the lower bound for the dilation factor is $1$; the upper bound for the dilation factor is computed as $\lfloor\frac{n}{k}\rfloor$, the greatest integer less than or equal to $\frac{n}{k}$, where $n$ is the feature size, and $k$ is the kernel size. By incrementing the value of dilation factor, the self-attention window size is widened, which means longer-range dependencies are learnt by the network. To capture both local and global blurred patterns, we design a cascaded structure where the dilation factor $\delta\in\left\{1,\lfloor\frac{n}{k}\rfloor\right\}$ takes the lower bound value and upper bound value alternately. $k$ is set to a constant number $7$ in DeblurDiNAT.
3.2 Channel Modulation Self Attention
Fig. 2 (b) and (c) show that CMSA contains two parallel units, DiNA and a local cross-channel learner (LCCL). As illustrated in Fig. 2 (d), an LCCL first transforms the normalized 2-D features into 1-D data by global average pooling (GAP); then, it applies a 1-D convolution with a fixed kernel size to the intermediate features along the channel dimension; finally, a softmax function is adopted to compute attention scores. Tensor manipulations are ignored in Fig. 2 (d). Element-wise multiplications are used to merge the outputs of the LCCL and DiNA. We claim that our design has two benefits. First, as there is no channel reduction in the LCCL, information loss along the channel dimension is avoided. Second, the parallel architecture optimizes the extra runtime caused by the LCCL. Extensive experiments show that our CMSA outperforms other similar variants, which are discussed in Sec. 4.3.
Given an input tensor $\mathbf{X}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$. The output of CMSA is computed as,
$\hat{\mathbf{X}}$ $=\mathrm{DiNA}(\mathbf{X})\odot\mathrm{LCCL}(\mathbf{X}),$ (4)
${\rm LCCL}(\mathbf{X})$ $=f^{-1}(\mathrm{Conv}_{1d}(f(\mathrm{GAP}_{2d}^{1\times 1}(\mathcal{L}(\mathbf{X}))))),$
where $\odot$ denotes element-wise multiplication; $\mathrm{DiNA}$ is explained in Eq. 1, Eq. 2 and Eq. 3; $f$ is a tensor manipulation function which squeezes and transposes a $C\times 1\times 1$ matrix, resulting in a $1\times C$ matrix; $\mathrm{Conv}_{1d}$ denotes a 1D convolution with a kernel size of $3$; ${\rm GAP}_{2d}^{1\times 1}$ indicates global average pooling, outputting a tensor of size $1\times 1$; $\mathcal{L}$ is layer normalization.
3.3 Divide and Multiply Feed-Forward Network
Figure 2: The structures of (a) Transformer block, (b) local CMSA with a dilation factor of $1$, (c) global CMSA with a dilation factor of $\frac{n}{k}$, (d) LCCL, and (e) DMFN. The kernel sizes used in 1D Conv and Depthwise Conv are set to $3$.
As shown in Fig. 2 (e), DMFN removes the activation function from regular feed-forward networks, motivated by the simple gate [3]. We claim that DeblurDiNAT is still a powerful non-linear model, because Leaky ReLU in the encoder path ensures that the extracted feature maps contain non-linearity, and feature fusion modules with GELU propagate features with non-linearity across the network. From a layer normalized feature $\mathbf{X}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, DMFN first applies point-wise convolutions to expand the feature channels by a factor $\gamma=2$. After that, $3\times 3$ depth-wise convolutions are utilized to aggregate spatial context, generating $\mathbf{X}_{0}\in\mathbb{R}^{\hat{H}\times\hat{W}\times 2\hat{C}}$. Next, the intermediate feature map $\mathbf{X}_{0}$ is equally divided along the channel dimension, yielding $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, so that their element-wise multiplication result is of shape $\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$. The overall process of DMFN with normalization is formulated as:
$\hat{\mathbf{X}}$ $=W_{p}^{0}\mathrm{DM}(\mathbf{X})+\mathbf{X},$ (5)
$\mathrm{DM}(\mathbf{X})$ $=W_{d}^{1}W_{p}^{1}\mathcal{L}(\mathbf{X})\odot W_{d}^{2}W_{p}^{2}\mathcal{L}(\mathbf{X}),$
where $W_{p}^{(\cdot)}$ and $W_{d}^{(\cdot)}$ denote the convolution with filter size of $1\times 1$ and $3\times 3$; $\mathrm{DM}$ represents our divide and multiply strategy.
3.4 Lightweight Dual-Stage Feature Fusion
Figure 3: Structures of (a) multi-scale LDFF, (b) same-scale LDFF, (c) ECR, (d) CFM, (e) multi-scale AFF [6], (f) same-scale simplified gated feature fusion (SGFF). Same-scale simplified feature fusion (SFF) is a regular convolutional layer.
We propose two variants of lightweight dual-stage feature fusion (LDFF), the multi-scale one and the same-scale one, as shown in Fig. 3 (a)-(b). As a preprocessing step, LDFF first resizes input features if they are from different encoder levels, and concatenate them along the channel dimension. The main body of LDFF is composed of two stages, efficient channel reduction (ECR) and complementary feature mixing (CFM). In ECR, a $1\times 1$ convolution is used to merge cross-channel contextual information, and a $3\times 3$ depth-wise convolution is for local content aggregation in the spatial domain. Then, CFM splits a normalized feature map into two complementary features by linear layers, passing through two parallel branches, resulting in a linear feature in the identity path and a non-linear one in the GELU gated path. The two complementary features are merged by element-wise multiplications and additional linear projection layers to refine image details. The two multi-scale LDFF blocks in DeblurDiNAT are formulated as:
$\mathrm{LDFF}_{1}^{out}=\mathrm{LDFF}_{1}(\mathrm{EB}_{1}^{out},(\mathrm{EB}_{2}^{out})^{\times 2},(\mathrm{EB}_{3}^{out})^{\times 4}),$ (6)
$\mathrm{LDFF}_{2}^{out}=\mathrm{LDFF}_{2}((\mathrm{EB}_{1}^{out})^{\times\frac{1}{2}},\mathrm{EB}_{2}^{out},(\mathrm{EB}_{3}^{out})^{\times 2}),$
[…]
4 Experiments and Analysis
4.1 Datasets and Implementation Details
The training dataset is GoPro [32] composed of 2103 blurred and clean image pairs for training and $1$ pairs for testing. We tested DeblurDiNAT on the two synthetic datasets GoPro [32] and HIDE [35], and evaluated deblurring methods on real-world datasets RealBlur-R [34] and RealBlur-J [34]. DeblurDiNAT was trained only on the GoPro [32] training dataset and directly applied to the GoPro [32], HIDE [35], RealBlur-R [34] and RealBlur-J [34] test datasets.
We propose DeblurDiNAT-S (Small) and DeblurDiNAT-L (Large), the two variants of our model. The number of Transformer blocks in DeblurDiNAT-S are $[4,6,8]$ at each level, attention heads are $[2,4,8]$, and the number of channels are $[64,128,256]$. The refinement stage contains 4 Transformer blocks. For DeblurDiNAT-L, one more residual block is appended to each encoder; the number of Transformer blocks increment to $[6,12,18]$. We trained both models using Adam optimizer with the initial learning rate of $2\times 10^{-4}$ which was gradually decayed to $10^{-7}$ by the cosine annealing strategy. The batch size was set to $8$. We augmented input data using random cropping, flipping and rotation. We pretrained DeblurDiNAT models for 4000 epochs with a patch size of $256\times 256$ and further trained them for another 2000 epochs with a patch size of $512\times 512$. DeblurDiNAT was trained on eight NVIDIA A100 GPUs.
4.2 Performance Comparison
Quantitative Analysis. We compare DeblurDiNAT with existing SOTA deblurring models, including CNNs [32, 26, 37, 12, 47, 6, 46] and Transformers [45, 41, 38, 24]. All models were only trained on the GoPro dataset and tested on the GoPro [32], HIDE [35], RealBlur-R [34] and RealBlur-J [34] datasets. The image quality metrics PSNR/SSIM, and inference time are tested with full resolution $1280\times 720$ images on a single NVDIA A100 GPU. The number of FLOPs and GPU consumption are calculated with an input tensor of size $256\times 256$ on an NVIDIA 3090 GPU. The GPU consumption is computed by the “torch.cuda.max memory allocated” function.
Tab. 1 reports that DeblurDiNAT-L obtains an average boost of $0.18$ dB over the previous SOTA approach FFTformer [24]. More specifically, Tab. 1 shows that DeblurDiNAT-L holds the second position in all the PSNR score rankings on synthetic datasets [32, 35] and real-world datasets [34]. It proves that our robust DeblurDiNAT-L demonstrates the strongest generalization ability. Tab. 2 compares the efficiency of our model to existing deblurring Transformers [45, 41, 38, 24]. It reports that DeblurDiNAT-L has $57\%$ less GPU memory consumption and runs 2.65×2.65$\times$2.65 × faster than FFTformer [24] which heavily involves complex matrix operations. The PSNR scores of DeblurDiNAT-L measured on the unknown RealBlur datasets [34] are higher than those of all models except Uformer-B [41]. Compared to Uformer-B [41], our DeblurDiNAT-L obtains an average gain of $0.23$ dB and runs 3.70×3.70$\times$3.70 × faster with $68\%$ fewer parameters. DeblurDiNAT-S outperforms nearly all models but FFTformer [24] on the synthetic GoPro dataset [32]. Compared to nearest competitors [6, 46, 45, 41, 38], DeblurDiNAT-S has $43.48\%$-$82.12\%$ fewer parameters. From quantitative comparisons Tab. 1 and Tab. 2, we argue that our DeblurDiNAT guarantees effectiveness, efficiency and robustness. Qualitative analysis (see Sec. 4.2) presents that our restored images are visually closest to the clean images.
Table 1: The average PSNR is calculated over the GoPro [32], HIDE [35], RealBlur-R [34] and RealBlur-J [34] datasets with equal weights to measure the generalization ability of each deblurring model. The best, second-best and third-best PSNR scores are highlighted, underlined and double-underlined.
Avg. GoPro HIDE RealBlur-R RealBlur-J
Method PSNR PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Nah et al. [32] 28.80 29.08 0.914 25.73 0.874 32.51 0.841 27.87 0.827
DeblurGAN-v2 [26] 30.03 29.55 0.934 26.61 0.875 35.26 0.944 28.70 0.866
SRN [37] 30.71 30.26 0.934 28.36 0.915 35.66 0.947 28.56 0.867
DSD [12] - 30.90 0.935 29.11 0.913 - - - -
DMPHN [47] 31.10 31.20 0.940 29.09 0.924 35.70 0.948 28.42 0.860
MIMO-Unet+ [6] 31.40 32.45 0.957 29.99 0.930 35.54 0.947 27.63 0.837
MPRNet [46] 32.08 32.66 0.959 30.96 0.939 35.99 0.952 28.70 0.873
Restormer [45] 32.32 32.92 0.961 31.22 0.942 36.19 0.957 28.96 0.879
Uformer-B [41] 32.31 33.06 0.967 30.90 0.953 36.19 0.956 29.09 0.886
Stripformer [38] 32.25 33.08 0.962 31.03 0.940 36.08 0.954 28.82 0.876
FFTformer [24] 32.36 34.21 0.969 31.62 0.946 35.87 0.953 27.75 0.853
DeblurDiNAT-S 32.18 33.15 0.963 30.92 0.939 35.93 0.953 28.73 0.874
DeblurDiNAT-L 32.54 33.63 0.967 31.47 0.944 36.09 0.955 28.98 0.885
Table 2: Model efficiency comparisons based on parameters, FLOPs, inference time, and GPU consumption.
Method Params (M) FLOPs (G) Time (ms) GPU Mem. (GB)
MIMO-Unet+ [6] 16.1 154.41 47.1 1.4
MPRNet [46] 20.1 777.01 46.3 5.8
Restormer [45] 26.1 140.99 21.00 11.8
Uformer-B [41] 50.9 85.72 676.1 3.4
Stripformer [38] 19.7 170.46 20.0 3.0
FFTformer [24] 16.6 - 483.8 19.0
DeblurDiNAT-S 9.1 99.46 24.3 5.2
DeblurDiNAT-L 16.1 152.61 182.8 8.2
Qualitative Analysis. The PSNR/SSIM metrics may not precisely measure the visual quality of deblurred images as perceived by human eyes. In this section, we present that the deblurred images produced by DeblurDiNAT-L stand out for their superior visual quality on both synthetic and real-world datasets [32, 35, 34]. All deblurring models are trained only on GoPro [32]. PSNR values are calculated based on the full-size images. Fig. 4 and Fig. 5 show that our method produces cleaner plate numbers and vivid facial details. Comparing a vast number of deblurred images using human eyes, we also notice that our DeblurDiNAT-L demonstrates a stronger generalization ability over existing models. For example, Fig. 6 shows that FFTformer [24] with the highest PSNR score tested on GoPro introduces unpleasant color changes when applied to the RealBlur-J dataset [34]; Uformer-B [41] with the best quantitative results on RealBlur datasets [34] might distort the appearance of texts according to our experiments. However, ours not only successfully preserves original picture colors but also clearly restores local details and structures of objects with real-world blur artifacts. Fig. 7 shows that DeblurDiNAT-L generates tidy texts and colorful graphic symbols from a blurred image in low-light environments, where the other networks fall short (feel free to zoom images in and out). From the above comparisons, we assert that DeblurDiNAT-L generates deblurred images with the best overall visual quality. More visual results are provided in supplementary materials.
(a) PSNR
Reference
(b) 22.44dB
Blurred
(c) 27.11dB
MIMO-Unet [6]
(d) 29.44dB
MPRNet [46]
(e) 28.76dB
Uformer-B [41]
(f) 29.56dB
Restormer [45]
(g) 30.16dB
Stripformer [38]
(h) 29.97dB
FFTformer [24]
(i) 30.21dB
DeblurDiNAT-S
(j) 30.43dB
DeblurDiNAT-L
Figure 4: Qualitative comparisons on the GoPro dataset [32].
(a) PSNR
Reference
(b) 22.07dB
Blurred
(c) 27.70dB
MIMO-Unet+ [6]
(d) 29.32dB
MPRNet [46]
(e) 29.11dB
Uformer-B [41]
(f) 29.37dB
Restormer [45]
(g) 29.39dB
Stripformer [38]
(h) 27.20dB
FFTformer [24]
(i) 29.39dB
DeblurDiNAT-S
(j) 30.39dB
DeblurDiNAT-L
Figure 5: Qualitative comparisons on the HIDE dataset [35].
(a) Reference
(b) Blurred
(c) DeblurGANv2 [26]
(d) MIMO-Unet+ [6]
(e) MPRNet [46]
(f) Uformer-B [41]
(g) Restormer [45]
(h) Stripformer [38]
(i) FFTformer [24]
(j) DeblurDiNAT-L
(k) Reference
(l) Blurred
(m) DeblurGANv2 [26]
(n) MIMO-Unet+ [6]
(o) MPRNet [46]
(p) Uformer-B [41]
(q) Restormer [45]
(r) Stripformer [38]
(s) FFTformer [24]
(t) DeblurDiNAT-L
Figure 6: Qualitative comparisons on the RealBlur-J dataset [34].
(a) Reference
(b) Blurred
(c) DeblurGANv2 [26]
(d) MIMO-Unet+ [6]
(e) MPRNet [46]
(f) Uformer-B [41]
(g) Restormer [45]
(h) Stripformer [38]
(i) FFTformer [24]
(j) DeblurDiNAT-L
Figure 7: Qualitative comparisons on the RealBlur-R dataset [34].
4.3 Ablation Study
In this section, we conduct ablation experiments to evaluate each component of DeblurDiNAT on the GoPro dataset [32]. We trained all models on two NVDIA A100 GPUs with a patch size of $128\times 128$ for 3000 epochs only.
Effect of ADFS.
Table 3: Investigation on Alternating Dilation Factor Structure.
Local or global PSNR/SSIM Time (ms) Params (M) FLOPs (G)
only local 31.73/0.951 19.6 9.082 99.41
only global 31.96/0.953 20.3 9.082 99.41
local+global 31.98///0.953 19.7 9.082 99.41
(a) Reference
(b) Blurred
(c) Only global
(d) Local+global
Figure 8: Effects of local attention on learning localized features.
Alternating Dilation Factor Structure (ADFS) is essentially a hybrid global-local attention strategy. As shown in Tab. 3, although DeblurDiNAT with hybrid attention and only global attention achieve similar metric scores, the introduction of local attention decreases the inference time by $2.96\%$ compared to the only global method. Visual comparions are demonstrated in Fig. 8, where local attention allows to capture region-specific blurred patterns and restore images with more visual details.
Table 4: Comparisons among combinations of different feature fusion methods. The component column is multi-scale fusion $+$ same-scale fusion.
Network Component PSNR Params (M) FLOPs (G)
Feature fusion AFF [6]$+$SFF $31.88$ 9.123 100.31
LDFF(SA$=$True)$+$SFF $31.88$ 9.096 99.71
LDFF(SA$=$True)$+$SGFF $31.89$ 9.079 99.28
LDFF$+$SGFF $31.89$ 8.995 97.13
LDFF$+$LDFF 31.96 9.082 99.41
Table 5: Influence of LCCL and DMFN. In the LCCL group, the positions of LCCL and impacts of different global pooling methods are investigated.
Network Component PSNR/SSIM FLOPs (G) Time (ms)
LCCL CMSA w/o LCCL $31.98/0.953$ 99.41 19.7
LCCL before SA$+$GAP $31.99/0.954$ 99.46 22.6
LCCL after SA$+$GAP $32.01/0.954$ 99.46 22.2
parallel$+$GMP $31.89/0.952$ 99.46 21.7
parallel$+$GAP 32.07///0.954 99.46 22.1
Feed-forward GDFN [45] $31.98/0.953$ 99.41 19.7
DMFN 32.05/0.954 99.41 19.4
Effect of LDFF. Several fusion units are illustrated in Fig. 3. In Tab. 4, replacing AFF [6] by LDFF (enhanced with SA) yields the same quality scores with $27$ thousands fewer parameters and $0.6$ G fewer FLOPs. Tab. 4 shows that the self-attention mechanism does not improve the metric scores but brings about additional computational cost in LDFF. It also presents that applying LDFF to both multi-scale and same-scale feature fusion provides an overall gain of $0.08$ dB and improved efficiency over the baseline. Moreover, removing the gating mechanism from LDFF results in a noticeable $0.01$dB performance drop in an architecture without any activation in feed-forward networks. The reason is that our non-linear LDFF with has a stronger ability to model and propagate complex data across DeblurDiNAT, of which the decoder path contains no activation.
Effect of LCCL. We investigate how different positions of cross-channel learner (LCCL) in CMSA and various pooling strategies used in LCCL affect deblurring performance. Tab. 5 shows that, our parallel design obtains a $0.08$ dB gain and $2.26\%$ speedup over the strategy placing LCCL before the self-attention blocks. The global average pooling (GAP) in LCCL leads to $0.18$ dB boost over the global max pooling (GMP). Compared to the self-attention without channel modulation, our proposed CMSA improves PSNR by $0.09$ dB.
Effect of DMFN. Experimental data in Tab. 5 demonstrates that our straightforward DMFN brings a performance improvement of $0.07$ dB PSNR compared to GDFN [45] with gating mechanisms. DMFN also offers a $1.5\%$ speedup. Both of them share similar numbers of FLOPs.
Analysis on Transformer Blocks. We compare CMSA with self-attention mechanisms used in Restormer [45], Stripformer [38] and FFTformer [24], three SOTA Transformers designed for image deblurring. For a fair comparison, We replace both the self-attention CMSA and feed-forward network DMFN in DeblurDiNAT with their [45, 38, 24] corresponding ones. As reported in Tab. 6, ours outperforms the other attention methods in terms of PSNR and GPU memory consumption. Compared to SA employed in FFTformer [24], ours achieves $0.04$ dB higher PSNR, 2.45×2.45$\times$2.45 × faster speed and $49\%$ less GPU memory consumption.
Table 6: Comparison experiments among different Transformer blocks.
self-attention PSNR Params (M) FLOPs (G) Time (ms) GPU Mem. (GB)
Restormer [45] 30.76 9.141 100.88 18.0 5.23
Stripformer [38] 31.21 8.776 97.19 64.2 5.90
FFTformer [24] 32.04 11.872 125.98 54.1 9.30
CMSA 32.08 9.082 99.46 22.1 4.73
5 Conclusion
In this work, we proposed a novel Transformer, called DeblurDiNAT, and instroduced four key designs to solve existing issues of deblurring Transformers. Firstly, an alternating dilation factor structure is utilized to recognize both local and global blurred patterns. Then, DeblurDiNAT uses a local cross-channel learner to modulate the self-attention outputs dynamically, so that cross-channel interactions can be learnt from complex image data. With the aim of efficient feature propagation, a divide and multiply feed-forward network replaces the conventional approaches. In addition, a lightweight dual-stage feature fusion method is created to introduce non-linearity and aggregate multi-scale or same-scale features. Comprehensive experiments demonstrate that our architecture outperforms SOTA models in regard to the efficiency and effectiveness balance, and ours exhibits the strongest generalization ability.
Acknowledgments. This work was supported by the Faster GPU cluster built by Texas A&M High Performance Research Computing (HPRC). We also thank NATTEN for the partial support of this work, which is an open-source deep learning library with efficient C++ and CUDA kernels.